OpenStudy (anonymous):

a - b = 4 2a -2b = 8 What is the answer please....

OpenStudy (callisto):

infinitely many solutions as the 2 equations are the same

OpenStudy (callisto):

2a -2b = 8 => a-b=4

OpenStudy (hoblos):

multiply the first equation by 2 2a -2b =8 2a -2b = 8 they are the same equation though then we have infinity of solutions